OpenStudy (anonymous):

find the equation of the line where the gradient and point on the line are given 1/2;(-2;3)

OpenStudy (amistre64):

given gradient: m and point: (Px, Py) y-Py = m(x-Px)

OpenStudy (anonymous):

use the general form y - y1 = m(x - x1) - as amis as indicated